### v1

@@ Abstract
Label scarcity problem is the main challenge that hinders the wide application of deep learning systems in automatic cardiovascular diseases (CVDs) detection using electrocardiography (ECG). Tuning pre-trained models alleviates this problem by transferring knowledge learned from large datasets to downstream small datasets. However, bottlenecks in computational efficiency and CVDs detection performance limit its clinical applications. It is difficult to improve the detection performance without significantly sacrificing model computational efficiency. Here, we propose a computation-efficient semi-supervised learning paradigm (FastECG) for robust and computation-efficient CVDs detection using ECG. It enables a robust adaptation of pre-trained models on downstream datasets with limited supervision and high computational efficiency. First, a random-deactivation technique is developed to achieve robust and fast low-rank adaptation of pre-trained weights. Subsequently, we propose a one-shot rank allocation module to determine the optimal ranks for the update matrices of the pre-trained weights. Finally, a lightweight semi-supervised learning pipeline is introduced to enhance model performance by leveraging labeled and unlabeled data with high computational efficiency. Extensive experiments on four downstream ECG datasets demonstrate that FastECG not only outperforms the state-of-the-art methods in multi-label CVDs detection but also consumes fewer GPU footprints, training time, and parameter storage space. As such, this paradigm provides an effective solution for achieving high computational efficiency and robust detection performance in the clinical applications of pre-trained models under limited supervision.
keywords:
Electrocardiograph, Pre-trained models, Semi-supervised learning, Computation-efficient, Cardiovascular diseases.
@@ 1 Introduction
Cardiovascular diseases have become the deadliest ’killer’ of human health in recent years[1]. As a non-invasive and low-cost tool, ECG provides a visual representation of the electrical activity of the heart and is widely used in the detection of various CVDs[2, 3]. Benefiting from recent progress in computing hardware, ECG-based deep learning systems have achieved notable success in automatic CVDs detection[4, 5, 6]. However, previous deep learning models required sufficient labeled samples to achieve satisfactory performance when trained on new application scenarios with unseen CVDs[7, 8]. However, collecting well-labeled ECG recordings requires physicians’ expertise and their laborious manual annotation, and therefore is expensive and time consuming in clinical practice[9, 10]. Additionally, constantly updating models to adapt to different downstream datasets often introduces heavy computational burdens[11]. As such, designing a fast and robust deep learning paradigm for ECG-based CVDs detection under limited supervision is highly in demand but remains challenging.
Recent advancements in pre-trained models (PMs) have enhanced the performance of deep learning models on the downstream datasets without large-scale labeled data[12, 13, 14]. A commonly used pipeline consists of pre-training over-parameterized backbone models on large-scale datasets, and then fine-tuning them on small downstream datasets in a supervised manner. In this process, the prior knowledge learned from the large datasets during pre-training is transferred to the downstream datasets, thereby improving the model performance on new application scenarios[15]. For instance, Kiyasseh et al.proposed a self-supervised method to pre-train a large model for ECG-based CVDs detection[16]. Their fine-tuned model demonstrated robust performance on downstream datasets with only 25% of labeled data. Subsequently, using 8 million ECG recordings, Akhil et al.pre-trained a vision-transformer model with self-supervised masked image modeling to improve the fine-tuning performance on downstream datasets[17].
@@
However, two bottlenecks still greatly limit the clinical application of CVDs detection systems based on PMs. (1) The bottleneck in CVDs detection performance. Fine-tuning of pre-trained models is currently conducted in a purely supervised manner. When the sample size of the labeled data is very scarce in the downstream datasets, model performance may drop due to over-fitting[18]. Fortunately, a large number of unlabeled data in the medical domain are relatively easy to collect. Semi-supervised learning (SSL) is able to extract sufficient information from the unlabeled data and outperform the supervised models trained with the same amount of labeled data[19, 8, 20, 21, 22]. For example, self-tuning integrates the exploration of unlabeled data and the knowledge transfer of PMs into a united framework, which significantly outperforms supervised fine-tuning on five downstream tasks[18]. Oliveira et al.conducts a comprehensive comparison analysis between SSL and supervised methods in ECG-based CVDs detection, demonstrating the superior performance of the SSL methods. Utilizing a knowledge distillation mechanism, MixTeacher delivers an effective SSL framework for atrial fibrillation detection under limited supervision, reducing the workload of ECG recordings annotation by 98%[9]. Despite their robust performance, existing SSL methods are mainly built on pseudo-label techniques and the weak-strong consistency training on unlabeled samples[7, 23, 8, 21, 24], which greatly increase the GPU memory footprint and computation time during model training. Furthermore, SSL methods using teacher-and-student models for knowledge distillation require two full copies of the corresponding PMs, thus increasing the model storage consumption. Therefore, it is necessary to address the bottleneck of computational efficiency during the performance enhancement of PMs using semi-supervised learning.
@@
(2) The bottleneck in computational efficiency. Fine-tuning and SSL methods both update all the parameters of PMs. Despite their effectiveness, both methods have the main drawback that the tuned models have the same number of parameters as PMs, thus being computationally inefficient as the PMs scale up. Specifically, they require saving the gradients of all the parameters in PMs and even the momentum parameters, resulting in large GPU memory footprints when training large PMs[25]. Additionally, each tuned model can be regarded as a full copy of the corresponding PMs, and therefore leading to high storage consumption when simultaneously tuned on multiple datasets [26]. To address this, parameter-efficient fine-tuning (PEFT) methods have been introduced to reduce the trainable parameters during model training [11, 27, 28]. For example, BitFit[27] only tunes the bias terms in pre-trained models while freezing the remaining parameters, greatly reducing the number of trainable parameters. Low-rank adaptation (LoRA) achieves the same goal by updating the pre-trained weights with low-rank decomposition matrices. Consequently, parameter-efficient methods provide an effective solution to achieve parameter-efficient semi-supervised learning by integrating with different SSL methods. Despite their parameter efficiency, a comprehensive study concluded that they could not achieve on-par performance with fine-tuning on some complex downstream tasks[29]. AdaLoRA and IncreLoRA further improve the performance of LoRA by allocating different ranks to different pre-trained weights based on their importance[26, 30]. However, the above performance improvement is achieved at the cost of increased training time for iterative importance estimation.
@@
Therefore, a dilemma is encountered: model performance improvement often comes at the expense of a large sacrifice of computational efficiency. Specifically, semi-supervised learning enhances CVDs detection performance, but at a significantly increased computational cost. Conversely, parameter-efficient methods that prioritize computational efficiency may compromise CVDs detection performance. Consequently, achieving a superior detection performance with high computation efficiency poses a great challenge to the clinical application of PMs in ECG-based CVDs detection.
@@
In this study, we propose a united paradigm capable of addressing the above two bottlenecks simultaneously. It is a computation-efficient semi-supervised learning paradigm (FastECG) for adapting pre-trained models on downstream datasets with high computational efficiency under limited supervision. Our method enables robust and low-cost detection of CVDs in clinical practice using ECG recordings. As shown in Fig.1, first, a base backbone is pre-trained on a large-scale 12-lead ECG dataset in a supervised manner, which provides a foundation model for downstream tasks. Second, a random-deactivation low-rank adaptation (RD-LoRA) method formulates a low-cost and robust pipeline for updating the pre-trained backbone on downstream datasets. Specifically, it stochastically activates or deactivates low-rank adaptation in each trainable layer of the backbone with a probability $p$. To reduce GPU memory footprint, the pre-trained weights in each layer are always frozen. Theoretical analysis indicates that the random deactivation operation integrates various sub-networks generated during model training, thus overcoming the performance bottleneck in tuning PMs. Additionally, deactivating low-rank adaptation in some layers reduces computation costs and speeds up the training process, especially when the backbone model size is large. Third, a one-shot rank allocation module allocates the optimal ranks for the low-rank matrices in each layer. In contrast to AdaLoRA[26] and IncreLoRA[30], the proposed method is able to determine the optimal ranks using only one gradient backward iteration, improving the adaptation performance at low computational costs.
@@
Additionally, a lightweight semi-supervised learning module is developed to leverage the abundant information within unlabeled data. This module uses unlabeled data to stabilize the statistics estimation in batch normalization layers, enhancing their generalization performance on unseen data distributions. Compared to the pseudo-labeling and the weak-strong consistency training methods[23, 8, 24], our lightweight module overcomes the label scarcity problem with significantly higher computational efficiency.
@@
Finally, extensive experiments on four downstream ECG datasets demonstrate the superior CVDs detection performance of the proposed FastECG against various state-of-the-art models under very limited supervision. Most importantly, our method only requires 66.5% training time, 70.7% GPU memory footprint, and 1.8%-5.8% trainable parameters of the state-of-the-art SSL methods. In conclusion, our proposed computation-efficient semi-supervised learning paradigm provides an effective solution to overcome the two bottlenecks that limit the clinical applications of PMs in ECG-based CVDs detection.
2 Results
Designing a robust and low-cost deep learning paradigm for CVDs detection under label-scarcity conditions remains a challenge in medical intelligence. The proposed FastECG addresses this challenge by introducing a novel computation-efficient semi-supervised learning paradigm (Fig.1), which integrates four important sub-modules. We conduct extensive and systematic experiments to evaluate its effectiveness and computational efficiency on various public databases. Our results indicate that FastECG not only outperforms state-of-the-art semi-supervised and parameter-efficient semi-supervised baselines in ECG-based CVD detection but also demonstrates satisfactory computational efficiency.

### v2

@@ Abstract
The label scarcity problem is the main challenge that hinders the wide application of deep learning systems in automatic cardiovascular diseases (CVDs) detection using electrocardiography (ECG). Tuning pre-trained models alleviates this problem by transferring knowledge learned from large datasets to downstream small datasets. However, bottlenecks in computational efficiency and detection performance limit its clinical applications. It is difficult to improve the detection performance without significantly sacrificing the computational efficiency during model training. Here, we propose a computation-efficient semi-supervised learning paradigm (CE-SSL) for robust and computation-efficient CVDs detection using ECG. It enables a robust adaptation of pre-trained models on downstream datasets with limited supervision and high computational efficiency. First, a random-deactivation technique is developed to achieve robust and fast low-rank adaptation of pre-trained weights. Subsequently, we propose a one-shot rank allocation module to determine the optimal ranks for the update matrices of the pre-trained weights. Finally, a lightweight semi-supervised learning pipeline is introduced to enhance model performance by leveraging labeled and unlabeled data with high computational efficiency. Extensive experiments on four downstream datasets demonstrate that CE-SSL not only outperforms the state-of-the-art methods in multi-label CVDs detection but also consumes fewer GPU footprints, training time, and parameter storage space. As such, this paradigm provides an effective solution for achieving high computational efficiency and robust detection performance in the clinical applications of pre-trained models under limited supervision. Code and Supplementary Materials are available at https://github.com/KAZABANA/CE-SSL
Index Terms:
Electrocardiograph; Semi-Supervised Learning; Cardiovascular Diseases; Computation-Efficient.
@@ 1 Introduction
Cardiovascular diseases have become the deadliest ’killer’ of human health in recent years[1]. As a non-invasive and low-cost tool, ECG provides a visual representation of the electrical activity of the heart and is widely used in the detection of various CVDs[2, 3]. Benefiting from recent progress in computing hardware, ECG-based deep learning systems have achieved notable success in automatic CVDs detection[4, 5, 6, 7]. However, previous deep learning models required sufficient labeled samples to achieve satisfactory performance when trained on new application scenarios with unseen CVDs[8, 9]. Unfortunately, collecting well-labeled ECG recordings requires physicians’ expertise and their laborious manual annotation, and therefore is expensive and time-consuming in clinical practice[10, 11]. Recent advancements in pre-trained models have enhanced the performance of deep learning models on the downstream datasets without large-scale labeled data[12, 13, 14]. A commonly used pipeline consists of pre-training over-parameterized backbone models on large-scale datasets and then fine-tuning them on small downstream datasets in a supervised manner. However, two bottlenecks still greatly limit the clinical application of CVDs detection systems based on pre-trained models under limited supervision.
@@
(1) The bottleneck in CVDs detection performance. Fine-tuning of pre-trained models is currently conducted in a purely supervised manner. When the labeled data is very scarce in the downstream datasets, model performance may drop due to over-fitting[15]. Fortunately, a large number of unlabeled data in the medical domain are relatively easy to collect. Semi-supervised learning (SSL) is able to extract sufficient information from the unlabeled data and outperform the supervised models trained with the same amount of labeled data[16, 9, 17, 18, 19]. For example, self-tuning integrates the exploration of unlabeled data and the knowledge transfer of pre-trained models into a united framework, which significantly outperforms supervised fine-tuning on five downstream tasks[15]. Despite their robust performance, existing SSL methods are mainly built on pseudo-label techniques and the weak-strong consistency training on unlabeled samples[8, 20, 9, 18, 21], which greatly increase the GPU memory footprint and computation time during model training. This drawback results in the bottleneck of computational efficiency during the performance enhancement of pre-trained models using semi-supervised learning.
@@
(2) The bottleneck in computational efficiency for parameter optimization. Nowadays, many studies have introduced large-scale foundation models to achieve better CVDs detection performance using ECG[22, 23, 24, 25, 26], greatly increasing the computation costs of modifying them for downstream applications. SSL methods and fine-tuning both update all the model parameters. Despite their effectiveness, both methods have the main drawback that they require saving the gradients of all the parameters and even the momentum parameters, resulting in large GPU memory footprints when tuning large pre-trained models[27]. Additionally, each tuned model can be regarded as a full copy of the original models, therefore leading to high storage consumption when simultaneously tuned on multiple datasets [28]. To address this, parameter-efficient fine-tuning (PEFT) methods have been introduced to reduce the trainable parameters during model training and thus decrease the computational costs during model training[29, 30, 31]. For example, Low-rank adaptation (LoRA) achieves this goal by updating the pre-trained weights with low-rank decomposition matrices. AdaLoRA and IncreLoRA overcome the performance bottleneck of LoRA by allocating different ranks to different pre-trained weights based on their importance[28, 32]. However, the above performance improvement is achieved at the cost of increased training time for iterative importance estimation.
@@
Therefore, a dilemma is encountered: model performance improvement often comes at the expense of a large sacrifice of computational efficiency during model training. Specifically, semi-supervised learning enhances CVDs detection performance under limited supervision but at significantly increased computational costs. Conversely, methods that prioritize computational efficiency may compromise model performance[33]. Consequently, achieving a superior detection performance with high computation efficiency poses a great challenge to the clinical application of pre-trained models in ECG-based CVDs detection. To the best of our knowledge, no prior study has designed and evaluated a framework to escape the dilemma.
@@
Here, we propose a united paradigm capable of addressing the above two bottlenecks simultaneously. It is a computation-efficient semi-supervised learning paradigm (CE-SSL) for adapting pre-trained models on downstream datasets with high computational efficiency under limited supervision. Our method enables robust and low-cost detection of CVDs in clinical practice using ECG recordings. As shown in Fig.1, first, a base backbone is pre-trained on a large-scale 12-lead ECG dataset in a supervised manner, which provides a foundation for downstream datasets. We also provide medium and large backbones for performance enhancement by increasing the backbone’s depth and width. Second, a random-deactivation low-rank adaptation (RD-LoRA) method formulates a low-cost and robust pipeline for updating the pre-trained backbone on downstream datasets. Specifically, it stochastically activates or deactivates low-rank adaptation in each trainable layer of the backbone with a probability $p$. To reduce GPU memory footprint, the pre-trained weights in each layer are always frozen. Theoretical analysis indicates that the random deactivation operation integrates various sub-networks generated during model training, thus overcoming the performance bottleneck in tuning pre-trained models. Additionally, deactivating low-rank adaptation in some layers reduces computation costs and speeds up the training process, especially when the backbone model size is large. Third, a one-shot rank allocation module allocates the optimal ranks for the low-rank matrices in each layer. In contrast to AdaLoRA[28] and IncreLoRA[32], the proposed method can determine the optimal ranks using only one gradient backward iteration, improving the adaptation performance at low computational costs.
@@
Additionally, a lightweight semi-supervised learning module is developed to leverage the abundant information within unlabeled data. This module uses unlabeled data to stabilize the statistics estimation process in batch normalization layers, enhancing their generalization performance on unseen data distributions. Compared to the pseudo-labeling and the weak-strong consistency training methods[20, 9, 21], our lightweight module alleviates the label scarcity problem with significantly higher computational efficiency.
@@
Finally, extensive experiments on four downstream datasets demonstrate the superior CVDs detection performance of the proposed CE-SSL against various state-of-the-art models under very limited supervision. Most importantly, our method only requires 66.5% training time, 70.7% GPU memory footprint, and 1.8%-5.8% trainable parameters of the state-of-the-art SSL methods. Furthermore, its computational costs can be minimized to adapt to resource-limited environments without a significant accuracy loss. In conclusion, our proposed computation-efficient semi-supervised learning paradigm provides an effective solution to overcome the two bottlenecks that limit the clinical applications of pre-trained models in ECG-based CVDs detection. We summarize the major contributions as follows:
  • Pre-trained backbones with various sizes are provided to serve as a foundation for ECG-based CVDs detection on downstream datasets.
  • A random deactivation low-rank adaptation method is proposed to update the backbones with high computational efficiency and robust performance. A one-shot rank allocation module is present to determine the optimal rank distribution during low-rank adaptation at minimal costs.
  • A lightweight semi-supervised method is proposed to leverage large-scale unlabeled data without greatly sacrificing the computational efficiency.
  • A computation-efficient semi-supervised framework for low-cost and accurate CVDs detection is proposed, which is the first one to escape the dilemma between model performance and computational efficiency.
2 Related Work
2.1 AI-Enabled CVDs Prediction using ECG
Benefiting from the development of deep learning, AI-enabled systems have shed light on automatic ECG screening and cardiovascular disease diagnosis[34, 4, 5, 35, 36, 37, 22, 23, 24]. Tracing the development of the systems, it can be observed that the prediction models they used are continuously scaling up. In the first stage, small-scale models demonstrated promising diagnosis performance in ECG analysis and CVDs detection. For example, Pourbabaee et al. designed a deep convolutional neural network to extract features from ECG signals and utilize standard classifiers for screening paroxysmal atrial fibrillation[34]. Hannun et al. proposed an end-to-end deep convolutional neural network to achieve automatic single-lead ECG screening[4]. The results demonstrated that the network achieved similar diagnosis performance compared with common cardiologists. In the second stage, pre-trained models with a prohibitive number of parameters were introduced, which demonstrated better transferability than previous networks. This advantage reduces their requirement for supervision information on downstream datasets. For instance, Vaid et al. pre-trained a large-scale vision transformer (HeartBEiT) on a huge ECG dataset and fine-tuned it on downstream datasets. The experiment results demonstrated the superiority of HeartBEiT in CVDs detection compared with traditional CNN architectures. In the current stage, many studies have proposed various kinds of foundation models for more advanced ECG screening and cardiac healthcare, inspired by their success in natural language processing[23, 24]. However, pre-trained models might experience a performance drop on downstream datasets when the labeled samples are very scarce there. Additionally, the computational costs of adapting them to various tasks significantly increase as their sizes scale up.
2.2 Semi-Supervised Learning for Performance Enhancement under Limited Supervision.
Semi-supervised learning offers an effective solution to address the label scarcity problem by leveraging unlabeled samples[15, 8, 9, 18, 21]. For example, Sohn et al. combined consistency regularization and pseudo-labeling to formulate a powerful algorithm (FixMatch)[9]. Extensive experiments demonstrate the superiority of FixMatch against the supervised baselines under the label scarcity condition. Subsequently, Zhang et al. proposed curriculum pseudo labeling (CPL) to flexibly adjust the thresholds for pseudo label selection, aiming at utilizing unlabeled data based on the model’s training progress[18]. Using a truncated Gaussian function, Chen et al. designed a soft threshold to weight unlabeled samples according to their prediction confidence, which achieved a balance between pseudo labels quality and quantity[21]. Compared with FixMatch, FlexMatch and SoftMatch both demonstrate better performance in various datasets. However, Wang et al. pointed out that the performance of semi-supervised models will be influenced by inaccurate pseudo labels, especially in large label space[15]. Hence, they proposed a self-tuning technique to explore the potential of the transfer of pre-trained models and a pseudo-label group contrast mechanism to increase the model’s tolerance to inaccurate labels. Experiments on five tasks demonstrated the superiority of the proposed framework against previous semi-supervised and supervised methods. In summary, massive unlabeled data and powerful pre-trained models led to the success of semi-supervised methods. However, high computation burdens are the side effects of leveraging them, greatly limiting their applications in resource-limited settings.
2.3 Parameter-Effcient Methods for Higher Computational Efficiency.
Parameter-efficient training has demonstrated great potential in decreasing the computational costs of fine-tuning pre-trained models[30, 27, 38, 28]. For example, Zaken et al. proposed BitFit to fine-tune the bias terms of the pre-trained models and freeze the other parameters, greatly reducing the computational costs. However, BitFit sacrifices the performance of the fine-tuned models because most of their parameters are not well adapted to downstream tasks. Hu et al. designed a low-rank adaptation method (LoRA) to inject trainable low-rank matrices into the transformer architecture, decreasing the performance gap between parameter-efficient methods and full fine-tuning[27]. However, Zhang et al. pointed out that LoRA ignored the varying importance of different pre-trained weights and allocated the same rank for all the trainable matrices, which led to suboptimal fine-tuning performance[28]. Consequently, they designed AdaLoRA to address this problem, which dynamically allocates different ranks to the low-rank matrices according to their importance during fine-tuning. During this process, the trainable parameters of the matrices with low importance are pruned. Different from AdaLoRA, IncreLoRA adaptively adds trainable parameters to the low-rank matrices with high importance[32]. As a non-pruning method, its performance is not limited by the preset parameter budget. Although IncreLoRA and AdaLoRA surpass LoRA in some scenarios, they result in high computation costs for weight importance estimation. Consequently, advancing fine-tuning performance without sacrificing computational efficiency remains challenging when designing parameter-efficient methods.